OpenStudy (anonymous):

find integration of (secx-1)^1/2

OpenStudy (callisto):

do you mean\[\int\limits\limits[ (\sec x)-1 ]^{(1/2)} dx ?\] or \[\int\limits\limits[ \sec( x-1) ]^{(1/2)} dx ?\]

OpenStudy (turingtest):

I see no better way than wolfram's http://www.wolframalpha.com/input/?i=+%28secx-1%29%5E1%2F2dx someone please show me a more elegant solution

OpenStudy (turingtest):

\(if\) it's the first one

OpenStudy (turingtest):

it must be\[\int\sqrt{\sec x-1}dx\]the other one is not really doable

OpenStudy (anonymous):

damn! I scribbled this integral with the solution on the margin of my notebook and now I can't find it!!!

OpenStudy (turingtest):

ok Fermat ;)

OpenStudy (anonymous):

its first one

OpenStudy (turingtest):

what kills me is how the solution from wolf is\[2\tanh^{-1}(\sqrt{\sec x+1})+C\]and then says "which for restricted values is equivalent to (a bunch of ugly crap)" still, I see no better way than wolf's here, and it's not that hard

OpenStudy (anonymous):

but whats the process i din't find any process in wolf

OpenStudy (turingtest):

click "show steps"

OpenStudy (callisto):

can anyone check it for me?

OpenStudy (anonymous):

i am not gettin step after taking (secx)-1=u

OpenStudy (anonymous):

hey callisto your attachment is not opening properly whould you plz write or draw it bro

OpenStudy (callisto):

i prefer sis to bro lol

OpenStudy (anonymous):

sorry callisto but both are wrong as in 1 we can't one variable out of integration and one in check at 6th step of your attachmenrt and the second one i didn't get it

OpenStudy (callisto):

the 6th step is...?

OpenStudy (turingtest):

you are correct in\[\int\sec xdx=\ln(|\sec x+\tan x|+C\]

OpenStudy (callisto):

@TuringTest and what's wrong with the other part? I need to learn from mistakes

OpenStudy (turingtest):

I'm not really sure what theorem you think youare applying in this step:\[\int(\sec x-1)^{1/2}dx={\frac32(\sec x-1)\over \int\sec xdx}+C\]that move really makes no sense to -why do you think you can divide by integral sec? -what happened to the 1/2 exponent on top? -where did 3/2 come from?

OpenStudy (callisto):

i mean to do something like ∫(ax+c)^n dx and that's sth i don't know how to deal with

OpenStudy (callisto):

so i tried the derivative way..

OpenStudy (turingtest):

I don't know what you mean the 'derivative way', but if you mean differentiation under the integral sign I think that technique is too advanced for you right now, and you have done it wrong. it looks like you tried to do this\[\int(\sec x-1)^{1/2}dx={\int(\sec x-1)^{1/2}\over \int\sec xdx}={\frac23(\sec x-1)^{3/2}\over \int\sec xdx}+C\]but the top cannot be integrated like that you also have no basis to divide the bottom by integral sec if you could just apply the power rule like that this problem would be a piece of cake, but you can't... do you see that\[\int(\sec x-1)^{1/2}dx\neq\frac32(\sec x-1)^{3/2}+C\]?

OpenStudy (callisto):

and what should i do?

OpenStudy (turingtest):

as I say in the above posts, this seems to be a pretty difficult integral The only way I can see to do it is wolframs way, though I am sure there is a better one out there somewhere

OpenStudy (callisto):

okay, thanks